OpenStudy (anonymous):

A boy is pushing a ring of mass 2kg and radius 0.5 m with a Stick the stick applies a Force of 2N on the ring and rolls it without slipping with accl of 0.3,/s^2 the coeffcient of friction b/w the ground and ring is large enough that rolling always occurs and coefficient of friction b/w stick and ring is(p/10) find p?

OpenStudy (anonymous):



OpenStudy (anonymous):

@CliffSedge @ghazi @ajprincess

OpenStudy (anonymous):

What is the direction of motion (force vector) of the stick?

OpenStudy (anonymous):

Also, is the acceleration linear 0.3 m/s^2 or angular 0.3 s^-2? It's a little ambiguous with that comma there.

OpenStudy (dls):

Find the Moment of Inertia of the ring I=MR^2=2*0.5*0.5=0.5

OpenStudy (dls):

a=alpha*r ==>0.3*0.5=0.15m/s^2

OpenStudy (dls):

Then find the torque using T=I*alpha then try using equations.. F-f=ma